# Interacting binaries on the Main Sequence as in-situ tracers of mass transfer efficiency and stability

### v1

@@ 2.4 Detailed binary evolution models
We calculate the cumulative mass transfer efficiency in detailed binary evolution models introduced in C. Wang et al. (2020, for the SMC) and K. Sen et al. (2022, for the LMC). At any timestep during the evolution of the binary, the cumulative mass transfer efficiency of the model is defined as in Eq. (1), where $M_{\rm d}$ and $M_{\rm a}$ denote the mass of the donor and accretor at that timestep, and $M_{\rm d,i}$ and $M_{\rm a,i}$ denote their initial masses at the Zero-Age Main Sequence. The contribution from each model is weighted by the initial mass function (E. E. Salpeter, 1955), the orbital period and mass ratio distribution (H. Sana et al., 2012, 2013), and the timestep during the Algol configuration (see Eq. (4) of K. Sen et al., 2022).
@@
While the detailed binary models include wind mass loss, we do not remove the contribution from the wind mass loss in calculating the mass transfer efficiency in the models. We discuss the effects of ignoring wind mass loss on our conclusions in Sect. 4.4. From our synthetic Algol population, we removed timesteps where mass transfer occurs from the initially less massive accretor to the donor, that is, when inverse mass transfer occurs. We also remove the timesteps where both binary components fill its Roche lobe (contact configuration, A. Menon et al., 2021; M. Fabry et al., 2022, 2023; J. Henneco et al., 2024; J. Vrancken et al., 2024)
@@
The stellar and binary physics assumptions in the models are described in detail in C. Wang et al. (2020); K. Sen et al. (2022). During a mass transfer phase, mass accretion onto the accretor is limited by the spin of the accretor star. Critically rotating accretors are assumed not to be able to accrete any more mass (N. Langer, 1998). The excess mass transferred is lost as winds from the accretor, carrying the specific orbitalangular momentum of the accretor (N. Langer et al., 2003; J. Petrovic et al., 2005). Mass transfer is assumed to be stable as long as the excess mass lost can be driven by the combined luminosity of both the binary components (see Eq. (2) of K. Sen et al., 2022). Strong tides (J. P. Zahn, 1977, implemented as in R. G. Detmers et al., 2008) in short-period ($\leq$5 d) systems can halt the accretor spin-up and lead to efficient mass accretion; longer period systems undergo inefficient mass transfer (K. Sen et al., 2022). This leads to an orbital period-dependent mass transfer efficiency in the models. When the combined luminosity of both stars is insufficient to drive the required excess mass loss, the binary is assumed to merge.
2.5 Mass transfer stability
@@
Mass transfer on the Main Sequence is not stable down to arbitrarily low values of initial mass ratios, owing to the increasingly divergent thermal timescales of the donor and the accretor (H. Ge et al., 2020). The mass transfer rate is typically set by the evolutionary timescale of the donor star (thermal initially and later nuclear), while the accretor can readjust its structure on its longer thermal timescale as compared to the donor. As such, most binaries with initial mass ratios ¡ 0.65 may merge during the thermal timescale mass transfer on the Main Sequence (S. E. de Mink et al., 2013; H. Ge et al., 2020).
@@
In the models of K. Sen et al. (2022, Fig. 1, green models), there exist two threshold values of initial mass ratios, such that 1) above the higher value $q_{\rm max}$ (bottom left of green region above the black shaded region), all binary models survive the thermal timescale mass transfer; 2) below the lower value $q_{\rm min}$ (top left of green region below the Case A boundary) of the initial mass ratio, all binary models are predicted to merge. The value of $q_{\rm max}$ and $q_{\rm min}$ are a function of the initial donor mass of the binary (K. Sen et al., 2023, Fig. 1). Between the two thresholds, the fate of the binary depends on its initial orbital period. Below the mass ratio $q_{\rm max}$, all the accretors spin up to critical rotation soon after the onset of thermal timescale mass transfer (W. Packet, 1981), and the mass transfer efficiency in such models is less than 10% (top panels of Fig. F2 of K. Sen et al., 2022).
@@
As a conservative estimate and for the sake of simplicity, we assume that all binaries may undergo arbitrarily efficient stable mass transfer up to an initial mass ratio given by $q_{\rm max}$ in the rotationally limited mass accretion scheme. Below $q_{\rm max}$, a higher mass transfer efficiency may be supported by the formation of a disk, where the coupling of the disk to the accretor can aid the accretion of mass while removing angular momentum from the accretor (B. Paczynski, 1991; R. Popham & R. Narayan, 1991). C. Schürmann & N. Langer (2024) investigate the response of non-rotating accretor stars for constant mass transfer efficiencies. We parametrise their results for the onset of L2 overflow, as a function of the initial donor mass and mass transfer efficiencies, to derive the minimum mass ratio for stable mass transfer in the presence of a disk (Fig. 8 of C. Schürmann & N. Langer, 2024).
@@
Table 3 lists the values of $M_{\rm ccd,i}$, $q_{\rm min}$, $q_{\rm max}$ and $P_{\rm orb,i,max}$ for different values of initial donor mass in the LMC and SMC. We linearly interpolate between the values and assume that the stellar and binary parameters for LMC metallicity are applicable for the Milky Way systems. The difference in values of $M_{\rm ccd,i}$ are within 0.1 $M_{\odot}$ between the LMC and SMC for the same initial donor mass. The values of $q_{\rm min}$, $q_{\rm max}$ for the two metallicities are separated by at most 0.05. The maximum initial orbital period $P_{\rm orb,i,max}$ is a significant function of metallicity, and we discuss in Sect. 3.2 that our conclusions on angular momentum loss are not strongly affected by metallicity.
@@
Figure 1: The parameter space of mass transfer efficiency $\varepsilon$ (bottom panels) and initial orbital period $P_{\rm orb,i}$ (top panels) as a function of the decreasing initial mass ratio $q_{\rm i}$ for LZ Cep (left panel, representing a typical case) and O026631 (right panel, representing an outlier system). The white region indicates the most likely values of mass transfer efficiency and specific angular momentum loss from the rotation-limited accretion model. Different colours correspond to curves of $P_{\rm orb,i}$ for alternative assumptions of $\gamma$=1 (red), 2 (green) and $M_{\rm d}/M_{\rm a}$ (purple, isotropic remission from the surface of the accretor). In the right panel, the purple star shows the initial orbital period for $\varepsilon=0$ (Eq. (7)). The black and blue hatched regions correspond to $\gamma\leq 0$ and $\gamma\geq 3$, respectively. The dark grey colour shows the region where all detailed binary evolution models merge. All detailed binary evolution models survive the thermal timescale mass transfer to the left of the light grey region (the highest value of $q_{\rm i}$ in the light grey region is the $q_{\rm max}$ limit). Models in the light grey region merge or survive depending on the binary orbital period; the shortest-period models merge, and vice versa (the lowest value of $q_{\rm i}$ in this region is the $q_{\rm min}$ limit). The orange dashed line shows the limiting mass transfer efficiency as a function of the initial mass ratio for which non-rotating models of C. Schürmann & N. Langer (2024) undergo L2 overflow. The orange shaded region, therefore, denotes the parameter space where an accretion disk-mediated mass transfer onto the accretor may not be stable. The upper limit to the top panels is given by $P_{\rm orb,i,max}$ calculated for each system.
3 Results
@@
In this section, we evaluate the possible range of mass transfer efficiency $\varepsilon$ and angular momentum loss factor $\gamma$ for each massive Algol binary in Table 1 and Table 2. We describe in detail our analysis for one representative binary (Sect. 3.1), discuss the outliers (Sect. 3.2), and then show the population properties in Sect. 3.3 and derive empirical constraints on mass transfer efficiency during the thermal timescale mass transfer phase on the Main Sequence.
3.1 Example analysis
@@
We describe our analysis in detail for the massive Algol in the Milky Way, LZ Cep (Fig. 1, left panel). The binary has a Roche-lobe filling donor star of mass 6.5 $M_{\odot}$ in a 3.07 d orbit with a 16 $M_{\odot}$ companion (L. Mahy et al., 2011). For the current donor mass $_{\rm d}=6.5MM_{\odot}$, we find the initial donor mass $_{\rm d,i}=13.46MM_{\odot}$ and maximum possible initial orbital period $P_{\rm orb,i,max}$ of 5.62 d (see Table 3). This implies that the donor has lost 6.96  $M_{\odot}$ during the fast Case A mass transfer phase. For the case of conservative mass transfer, the accretor could have accreted all the mass lost by the donor (neglecting wind mass loss), and the minimum initial mass of the accretor could be 9.04 $M_{\odot}$. Since the initial mass of the accretor cannot be higher than that of the donor, the maximum initial accretor mass can be 13.46 $M_{\odot}$. Hence, the initial mass of the accretor can range between 9.04-13.46 $M_{\odot}$.
@@
We can derive the plausible range of specific angular momentum loss factor $\gamma$ that can lead to the current orbital period $P_{\rm orb}=3.07$ d of the binary, for each value of $\varepsilon\neq 1$. The black hatched region is excluded because $\gamma<0$ implies the binary gains angular momentum on mass loss. We deem the blue shaded region unlikely because the system has to lose a high amount of angular momentum ($\gamma>3$, M. Nuijten & G. Nelemans, 2025). To the left of the grey regions, we find the plausible range of the initial orbital period $P_{\rm orb,i}$, and in turn $\gamma$. We set the lower limit to the initial orbital period to 1 d as most binaries having shorter orbital periods are expected to enter into a nuclear timescale contact phase shortly after the onset of mass transfer on the Main Sequence (A. Menon et al., 2021; M. Abdul-Masih et al., 2021, 2022; M. Fabry et al., 2022, 2023; M. J. Rickard & D. Pauli, 2023; J. Henneco et al., 2024; J. Vandersnickt & M. Fabry, 2025; M. Gull et al., 2025). The allowed range of $\gamma$ for this system is $\sim 0.5-3.0$. We also note that the binary is less likely to have originated from the most extended initial orbital periods due to considerations of angular momentum loss (blue hatched regions have $\gamma>3$).
@@ 3.2 Outliers
Detailed binary evolution models predict that the mass transfer leads to mergers at increasing initial mass ratios for decreasing initial donor masses (Table 3). We find that our analytical approach leads to no or highly inefficient solutions in most of the least massive Algols, both in the Milky Way and the SMC (see the top few rows of Fig. 5 and Fig. 6). The least massive Algols have the strongest constraints from the mass transfer stability criterion ($q_{\rm min}\sim 0.75$). Hence, we conclude that the least massive Algols give evidence that mass transfer must be stable up to $q_{\rm min}\sim 0.6$ at initial donor masses $_{\rm d,i}\approx 10-15MM_{\odot}$. This implies that the difference in stability of mass transfer when using detailed models compared to when using semi-analytic fits from J. R. Hurley et al. (2002) may impact rates of astrophysical events of interest, including neutron star X-ray binaries and binary neutron star mergers (X. T. Xu et al., 2025, see also J. J. Eldridge et al., 2017; P. Agrawal et al., 2023; T. Fragos et al., 2023; A. Chattaraj et al., 2025).
@@
To obtain a non-zero solution to the mass transfer efficiency, our detailed binary models will require a mechanism to remove angular momentum from the accretor while continually accreting mass, in the form of a decretion disk or magnetic braking. The orange dashed line shows the minimum mass ratio as a function of mass transfer efficiency such that non-rotating accretors do not undergo L2 overflow (C. Schürmann & N. Langer, 2024). The intersection of the orange dashed line with the black line represents the maximum mass transfer efficiency that the system could achieve in the presence of a disk to siphon angular momentum out of the accretor. For O026631, the maximum mass transfer efficiency $\varepsilon_{\rm max}^{\rm disk}\sim$ 0.45. We also note that rotation, however, can reach critical values in the accretor’s outer layers and have a large effect on the radius; therefore, we consider the models of C. Schürmann & N. Langer (2024) conservative in the sense that nature may reach L2 overflow at much higher initial mass ratios than inferred from non-rotating models. Moreover, it has been recently shown that L2 mass loss can occur when the decretion disk itself extends beyond the L2 point, significantly shrinking the orbit and leading to possible merger scenarios well before the non-rotating limit above (W. Lu et al., 2023).
3.3 Population properties
In this section, we derive the possible range of mass transfer efficiencies if each of the observed systems had undergone rotationally limited mass transfer only (Sect. 3.3.1), or for the case of a disk-mediated mass transfer that can siphon angular momentum from the accretor to the orbit while the accretor continually accretes mass (Sect. 3.3.2). We calculate the maximum value of mass transfer efficiency possible in the rotationally limited mass accretion case for each system at $q_{\rm max}$, and tabulate our results as $\varepsilon_{\rm qmax}$ in Table 1 and Table 2. We also show the minimum and maximum values of the angular momentum loss factor, tabulated as $\gamma_{\rm min,qmax}$ and $\gamma_{\rm max,qmax}$ in Table 1 and Table 2. We also show the maximum value of mass transfer efficiency possible for a disk-supported mass transfer $\varepsilon_{\max}^{\rm disk}$ in Table 1 and Table 2.
For about $\sim$50% massive Algols in the LMC (e.g. HV 2241, VFTS 652, SC1 105) and Milky Way (e.g. XX Cas, AB Cru), we find that non-conservative mass transfer ($\varepsilon=0$) is ruled out because the allowed range of initial accretor masses has to be lower than the initial donor mass (mass budget). In these systems, the current accretor masses are higher than the calculated initial donor masses, such that the accretors must have accreted mass from the donor to reach the current configuration. However, only two out of 29 systems in the SMC necessarily require such non-zero solutions to mass transfer efficiency.

### v2

@@ II.4 Detailed binary evolution models
We calculate the cumulative mass transfer efficiency in detailed binary evolution models introduced in Wang et al. (2020, for the SMC) and Marchant (2017, for the LMC); Sen et al. (2022, for the LMC). At any timestep during the evolution of the binary, the cumulative mass transfer efficiency of the model is defined as in Eq. (1), where $M_{\rm d}$ and $M_{\rm a}$ denote the mass of the donor and accretor at that timestep, and $M_{\rm d,i}$ and $M_{\rm a,i}$ denote their initial masses at the Zero-Age Main Sequence. The contribution from each model is weighted by the initial mass function (Salpeter, 1955), the orbital period and mass ratio distribution (Sana et al., 2012, 2013), and the timestep during the Algol configuration (see Eq. (4) of Sen et al., 2022).
@@
While the detailed binary models include wind mass loss, we do not remove the contribution from the wind mass loss in calculating the mass transfer efficiency in the models. We discuss the effects of ignoring wind mass loss on our conclusions in Sect. IV.4. From our synthetic Algol population, we removed timesteps where mass transfer occurs from the initially less massive accretor to the donor, that is, when inverse mass transfer occurs. We also remove the timesteps where both binary components fill its Roche lobe (contact configuration, Menon et al., 2021; Fabry et al., 2022, 2023; Henneco et al., 2024; Vrancken et al., 2024)
@@
The stellar and binary physics assumptions in the models are described in detail in Wang et al. (2020); Sen et al. (2022). During a mass transfer phase, mass accretion onto the accretor is limited by the spin of the accretor star. Critically rotating accretors are assumed not to be able to accrete any more mass (Langer, 1998). The excess mass transferred is lost as winds from the accretor, carrying the specific orbital angular momentum of the accretor (Langer et al., 2003; Petrovic et al., 2005). Mass transfer is assumed to be stable as long as the excess mass lost can be driven by the combined luminosity of both the binary components (see Eq. (2) of Sen et al., 2022). Strong tides (Zahn, 1977, implemented as in Detmers et al., 2008) in short-period ($\leq$5 d) systems can halt the accretor spin-up and lead to efficient mass accretion; longer period systems undergo inefficient mass transfer (Sen et al., 2022). This leads to an orbital period-dependent mass transfer efficiency in the models (see Fig. F2 of Sen et al., 2022 for quantitative numbers). When the combined luminosity of both stars is insufficient to drive the required excess mass loss, the binary is assumed to merge.
II.5 Mass transfer stability
@@
Mass transfer on the Main Sequence is not stable down to arbitrarily low values of initial mass ratios, owing to the increasingly divergent thermal timescales of the donor and the accretor (Ge et al., 2020). The mass transfer rate is typically set by the evolutionary timescale of the donor star (thermal initially and later nuclear), while the accretor can readjust its structure on its longer thermal timescale as compared to the donor. As such, most binaries with initial mass ratios << 0.65 may merge during the thermal timescale mass transfer on the Main Sequence (de Mink et al., 2013; Ge et al., 2020).
@@
In the models of Sen et al. (2022, Fig. 1, green models), there exist two threshold values of initial mass ratios, such that 1) above the higher value $q_{\rm max}$ (bottom left of green region above the black shaded region), all binary models survive the thermal timescale mass transfer; 2) below the lower value $q_{\rm min}$ (top left of green region below the Case A boundary) of the initial mass ratio, all binary models are predicted to merge. The value of $q_{\rm max}$ and $q_{\rm min}$ are a function of the initial donor mass of the binary (Sen et al., 2023, Fig. 1). Between the two thresholds, the fate of the binary depends on its initial orbital period. Below the mass ratio $q_{\rm max}$, all the accretors spin up to critical rotation soon after the onset of thermal timescale mass transfer (Packet, 1981), and the mass transfer efficiency in such models is less than 10% (top panels of Fig. F2 of Sen et al., 2022).
@@
As a conservative estimate and for the sake of simplicity, we assume that all binaries may undergo arbitrarily efficient stable mass transfer up to an initial mass ratio given by $q_{\rm max}$ in the rotationally limited mass accretion scheme. Below $q_{\rm max}$, a higher mass transfer efficiency may be supported by the formation of a disk, where the coupling of the disk to the accretor can aid the accretion of mass while removing angular momentum from the accretor (Paczynski, 1991; Popham and Narayan, 1991). Schürmann and Langer (2024) investigate the response of non-rotating accretor stars for constant mass transfer efficiencies. We parametrise their results for the onset of L2 overflow, as a function of the initial donor mass and mass transfer efficiencies, to derive the minimum mass ratio for stable mass transfer in the presence of a disk (Fig. 8 of Schürmann and Langer, 2024).
@@
Table 4 lists the values of $M_{\rm ccd,i}$, $q_{\rm min}$, $q_{\rm max}$ and $P_{\rm orb,i,max}$, as described above, for different values of initial donor mass in the LMC and SMC. We linearly interpolate between the values and assume that the stellar and binary parameters for LMC metallicity are applicable for the Milky Way systems. The difference in values of $M_{\rm ccd,i}$ are within 0.1 $M_{\odot}$ between the LMC and SMC for the same initial donor mass. The values of $q_{\rm min}$, $q_{\rm max}$ for the two metallicities are separated by at most 0.05. The maximum initial orbital period $P_{\rm orb,i,max}$ is a significant function of metallicity, and we discuss in Sect. III.2 that our conclusions on angular momentum loss are not strongly affected by metallicity.
@@
Figure 1: The parameter space of mass transfer efficiency $\varepsilon$ (bottom panels) and initial orbital period $P_{\rm orb,i}$ (top panels) as a function of the decreasing initial mass ratio $q_{\rm i}$ for LZ Cep (left panel, representing a typical case) and O026631 (right panel, representing an outlier system). The white regions in both panels indicates the most likely values of mass transfer efficiency and specific angular momentum loss from the rotation-limited accretion model. Different colours correspond to curves of $P_{\rm orb,i}$ for alternative assumptions of $\gamma$=1 (red), 2 (green) and $M_{\rm d}/M_{\rm a}$ (purple, isotropic remission from the surface of the accretor). In the right panel, the purple star shows the initial orbital period for $\varepsilon=0$ (Eq. (7)). The black and blue hatched regions correspond to $\gamma\leq 0$ and $\gamma\geq 3$, respectively. The dark grey colour shows the region where all detailed binary evolution models merge. All detailed binary evolution models survive the thermal timescale mass transfer to the left of the light grey region (the highest value of $q_{\rm i}$ in the light grey region is the $q_{\rm max}$ limit). Models in the light grey region merge or survive depending on the binary orbital period; the shortest-period models merge, and vice versa (the lowest value of $q_{\rm i}$ in this region is the $q_{\rm min}$ limit). The orange dashed line shows the limiting mass transfer efficiency as a function of the initial mass ratio for which non-rotating models of Schürmann and Langer (2024) undergo L2 overflow. The orange shaded region, therefore, denotes the parameter space where an accretion disk-mediated mass transfer onto the accretor may not be stable. The upper limit to the top panels is given by $P_{\rm orb,i,max}$ calculated for each system.
III Results
@@
In this section, we evaluate the possible range of mass transfer efficiency $\varepsilon$ and angular momentum loss factor $\gamma$ for each massive Algol binary in Table 2 and Table 3. We describe in detail our analysis for one representative binary (Sect. III.1), discuss the outliers (Sect. III.2), and then show the population properties in Sect. III.3 and derive empirical constraints on mass transfer efficiency during the thermal timescale mass transfer phase on the Main Sequence.
III.1 Example analysis
@@
We describe our analysis in detail for the massive Algol in the Milky Way, LZ Cep (Fig. 1, left panel). The binary has a Roche-lobe filling donor star of mass 6.5 $M_{\odot}$ in a 3.07 d orbit with a 16 $M_{\odot}$ companion (Mahy et al., 2011). For the current donor mass $_{\rm d}=6.5MM_{\odot}$, we find the initial donor mass $_{\rm d,i}=13.46MM_{\odot}$ and maximum possible initial orbital period $P_{\rm orb,i,max}$ of 5.62 d (see Table 4). This implies that the donor has lost 6.96  $M_{\odot}$ during the fast Case A mass transfer phase. For the case of conservative mass transfer, the accretor could have accreted all the mass lost by the donor (neglecting wind mass loss), and the minimum initial mass of the accretor could be 9.04 $M_{\odot}$. Since the initial mass of the accretor cannot be higher than that of the donor, the maximum initial accretor mass can be 13.46 $M_{\odot}$. Hence, the initial mass of the accretor can range between 9.04-13.46 $M_{\odot}$.
@@
We can derive the plausible range of specific angular momentum loss factor $\gamma$ that can lead to the current orbital period $P_{\rm orb}=3.07$ d of the binary, for each value of $\varepsilon\neq 1$. The black hatched region is excluded because $\gamma<0$ implies the binary gains angular momentum on mass loss. We deem the blue shaded region unlikely because the system has to lose a high amount of angular momentum ($\gamma>3$, Nuijten and Nelemans, 2025). To the left of the grey regions, we find the plausible range of the initial orbital period $P_{\rm orb,i}$, and in turn $\gamma$ [^1] The analytical equations for gamma (Eq. 6 & 7) are valid only when epsilon is constant throughout the mass transfer phase. Since the mass transfer efficiency in the detailed binary models is a function of time, the analytical line showing $j_{\rm loss}=j_{\rm acc-orb}$ does not directly correspond to the detailed binary evolution model’s specific angular momentum loss.. We set the lower limit to the initial orbital period to 1 d as most binaries having shorter orbital periods are expected to enter into a nuclear timescale contact phase shortly after the onset of mass transfer on the Main Sequence (Menon et al., 2021; Abdul-Masih et al., 2021, 2022; Fabry et al., 2022, 2023; Rickard and Pauli, 2023; Henneco et al., 2024; Vandersnickt and Fabry, 2025; Gull et al., 2025). The allowed range of $\gamma$ for this system is $\sim 0.5-3.0$. We also note that the binary is less likely to have originated from the most extended initial orbital periods due to considerations of angular momentum loss (blue hatched regions have $\gamma>3$).
@@ III.2 Outliers
Detailed binary evolution models predict that the mass transfer leads to mergers at increasing initial mass ratios for decreasing initial donor masses (Table 4). We find that our analytical approach leads to no or highly inefficient solutions in most of the least massive Algols, both in the Milky Way and the SMC (see the top few rows of Fig. 5 and Fig. 6). The least massive Algols have the strongest constraints from the mass transfer stability criterion ($q_{\rm min}\sim 0.75$). Hence, we conclude that the least massive Algols give evidence that mass transfer must be stable up to $q_{\rm min}\sim 0.6$ at initial donor masses $_{\rm d,i}\approx 10-15MM_{\odot}$. This implies that the difference in stability of mass transfer when using detailed models compared to when using semi-analytic fits from Hurley et al. (2002) may impact rates of astrophysical events of interest, including neutron star X-ray binaries and binary neutron star mergers (Xu et al., 2025, see also Eldridge et al., 2017; Agrawal et al., 2023; Fragos et al., 2023; Chattaraj et al., 2025).
@@
To obtain a non-zero solution to the mass transfer efficiency, our detailed binary models will require a mechanism to remove angular momentum from the accretor while continually accreting mass, in the form of a decretion disk or magnetic braking. The orange dashed line shows the minimum mass ratio as a function of mass transfer efficiency such that non-rotating accretors do not undergo L2 overflow (Schürmann and Langer, 2024). The intersection of the orange dashed line with the black line represents the maximum mass transfer efficiency that the system could achieve in the presence of a disk to siphon angular momentum out of the accretor. For O026631, the maximum mass transfer efficiency $\varepsilon_{\rm max}^{\rm disk}\sim$ 0.45. We also note that rotation, however, can reach critical values in the accretor’s outer layers and have a large effect on the radius; therefore, we consider the models of Schürmann and Langer (2024) conservative in the sense that nature may reach L2 overflow at much higher initial mass ratios than inferred from non-rotating models. Moreover, it has been recently shown that L2 mass loss can occur when the decretion disk itself extends beyond the L2 point, significantly shrinking the orbit and leading to possible merger scenarios well before the non-rotating limit above (Lu et al., 2023).
III.3 Population properties
In this section, we derive the possible range of mass transfer efficiencies if each of the observed systems had undergone rotationally limited mass transfer only (Sect. III.3.1), or for the case of a disk-mediated mass transfer that can siphon angular momentum from the accretor to the orbit while the accretor continually accretes mass (Sect. III.3.2). We calculate the maximum value of mass transfer efficiency possible in the rotationally limited mass accretion case for each system at $q_{\rm max}$, and tabulate our results as $\varepsilon_{\rm qmax}$ in Table 2 and Table 3. We also show the minimum and maximum values of the angular momentum loss factor, tabulated as $\gamma_{\rm min,qmax}$ and $\gamma_{\rm max,qmax}$ in Table 2 and Table 3. We also show the maximum value of mass transfer efficiency possible for a disk-supported mass transfer $\varepsilon_{\max}^{\rm disk}$ in Table 2 and Table 3.
For about $\sim$50% of the massive Algols in the LMC (e.g. HV 2241, VFTS 652, SC1 105) and Milky Way (e.g. XX Cas, AB Cru), we find that non-conservative mass transfer ($\varepsilon=0$) is ruled out because the allowed range of initial accretor masses has to be lower than the initial donor mass (mass budget). In these systems, the current accretor masses are higher than the calculated initial donor masses, such that the accretors must have accreted mass from the donor to reach the current configuration. However, only two out of 29 systems in the SMC necessarily require such non-zero solutions to mass transfer efficiency.